Reports 1-1 of 1 Clear search Modify search
AdV-DET (Commissioning)
gouaty, masserot, was - 17:14 Thursday 16 October 2025 (67961) Print this report
Comment to Glitch on the horizon visible in B1_DC_NULL (67950)

The SDB2_B1_DC_NULL_30_300_RMS channel is a useful trigger for this king of event .  At the begining of the 2025-10-15-13h LN3 lock segment there was several events with high SDB2_B1_DC_NULL_30_300_RMS values .

The one reported refer to the  2025-10-15-14h11m events (red rectangle) . 

As trial, the SDB2_B1_PD{1,2}_{Audio,DC} channels were reprocessed  to compute   SDB2_B1_PD{1,2}_Blended channels (plot and it zoom) :

  • data segment of  600s between 2025-10-15-14h03m02-UTC GPS144457200  and  2025-20-15-14h13m02-UTC GPS144457800
  • the raw  channels are  with the _ref suffix , the reprocessed ones  without 
  • There is no glitch on the reprocessed channels
    • this excludes an issue due to different filter computing method use to build the DC and Audio channels 
    • But it suggests a possible sample discrepancy between the B1_PD1 and B1_PD2 channels (Audio,DC) that are provided by 2 different photodiodes mezzanines
      • as a trial  on the SDB2_rtpc, the Tpro cycle could be delayed by a few us (5 us) to wait for the full data availability
Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×